OpenStudy (iesmee):

help solve equation: 3(x+4)=3(8-2x)

OpenStudy (anonymous):



OpenStudy (iesmee):

sorta makes sense, thank you

OpenStudy (anonymous):

3(x+4)=3(8-2x) x+4=8-2x 3x=4 x=4/3

OpenStudy (anonymous):

oh...I did not switch it